A tank holding milk has an elliptical manhole in it. The top of the manhole is 1.25m below the surface of the milk. The manhole is 750mm wide and 800mm high. Using the analytical method calculate the total force on the manhole and where it acts. The milk has an s.g of 1.026


1
Expert's answer
2021-06-24T15:08:02-0400

"Volume = L*w*h = \frac{800}{1000}*\frac{750}{1000}*1.25=0.75 m^3"

But we know that "W= mg = specific \space * gravity = 1.026 *9.8 = 10.0548 \frac{N}{m^3}"

Form the above information, "1 m^3 = 10.0548 N \implies 0.75 m^3 = \frac{10.0548 * 0.75 }{1}= 7.5411 N"